Commodities x and z are gross substitutes and so are x and y. Commodities x and z are gross substitutes but x and y are complements c. d. ng response question: Discuss the meaning of elasticity and the various types. What determines the price etasticity of demand for a certain good? Who is likely to find this information useful? Assume that the market demand for barley is given by: 1. a. b. Q-1,900-4P0.1M+2P Where Q is the quantity of barley demanded, P, is the price of barley, M is income (say per capita income of consumers) and P, is the price of wheat. The prices of wheat and barley are each 200 (say Es per tonne) and M is 1,000. The slopes for barley demand, wheat demand and income are -4, 2 and 0.1 respectively. Calculate the own price elasticity of demand, the income elasticity of demand and the cross-price elasticity of the demand for barley with respect to the price of wheat. Calculate and illustrate graphically the impact on welfare of a specific tax of 37.5p per unit to be paid by suppliers when Q 30 -4P and Qs-:-6 + 8R How do the welfare implications change if the tax is paid by consumers instead of suppliers? eu c. Interpret the following elasticities for petrol: Demand elasticity: -0.39 Income elasticity: 1.2 Supply elasticity: 0.7 a. Do there appear to be good substitutes for petrol in the preferences 2. of buyers? 45
Here isn’t the income elasticity meant to be 0.1 x 100/1600 because it’s slope x income /quantity
??
Answer #1

Income elasticity = Slope of demand curve with respect to income * Income / Quantity

Here,

Pw =Pb =200

Income = M = 1000

Slope of demand with respect to income = 0.1

At Pb = Pw = 200

Q = 1900 - 4(200) + 0.1*(1000) + 2(200)

Q = 1600

Income elasticity = 0.1*1000 / 1600 = 0.0625